OpenStudy (secret-ninja):

The numbers 20, 21, 29. form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? 21, 22, 30 40, 42, 58 60, 66, 87 70, 84, 116

OpenStudy (kirbykirby):

3 numbers\(a,b,c\) will be a pythagorean triple if it satisfies the equation \(a^2+b^2=c^2\), where \(c\) is the largest number out of the 3 numbers

OpenStudy (shiraz14):

@kirbykirby: That's correct.

OpenStudy (kirbykirby):

So it's really just a matter of testing all of your options, and seeing if that equation is true! :)

OpenStudy (secret-ninja):

so which one is it?

OpenStudy (kirbykirby):

Well just test them out... for example, the first one... you just try 21^2 + 22^2, and verify if it's equal to 30^2

OpenStudy (kirbykirby):

\(21^2+22^2=925\), now just verify if 925 is equal to \(30^2\)... if it's not equal, then those numbers are not a Pyhagorean triple

OpenStudy (secret-ninja):

im not very good at this. too bad no one ever gives the answer -.- :(

OpenStudy (kirbykirby):

Because we want you to understand the problem. If we just give the answer, you will not know how to do them on a test :(

OpenStudy (secret-ninja):

in school they only teach us what we need in test, not in real life. i dont think i need math to be a dog sitter lol

OpenStudy (secret-ninja):

at least not complicated math

OpenStudy (kirbykirby):

Well you should know basic math for that. And rally all you are doing here is inputting numbers into a calculator... which could be something you do as a dog sitter when calculating your pay

OpenStudy (secret-ninja):

ok then

OpenStudy (kirbykirby):

I don't want to come off as being mean :( But you should put a bit of effort into finding the solution rather than wait for people to give it to you. If you are struggling, then we can give you more assistance

OpenStudy (secret-ninja):

i just never even learnd this before. i dont want to learn it :(